During an election campaign a candidate gave away 5000 little advertisements. The pens she gave away cost $.23 cents each and the little bumper stickers cost $.14 cents each. He spent a total of $970.00. How many pens did he give away?

thanx

Found 2 solutions by fractalier, atif.muhammad:
Answer by fractalier(6550)

Let p be the number of pens and b be the number of bumper stickers.
Their costs are then .23p and .14b, respectively...so we have
p + b = 5000
.23p + .14b = 970
We can combine these by a linear sum by multiplying the first equation by 14 and the second by 100...then we subtract the equations...we get
14p + 14b = 70000
-(23p + 14b = 97000)
and get
-9p = - 27000
p = 3000
Thus he/she gave away 3000 pens and 2000 bumper stickers.

Answer by atif.muhammad(135)

Let's say he gave away x pens and y stickers.

He gave away 5000 little advertisements.

Therefore, x + y = 5000

The pens each cost 23 cents and the stickers cost 14 cents and he spent 970 dollars,

Therefore, 0.23x + 0.14y = 970

We now have simultaneous equations.

 x + y = 5000 (1)
0.23x + 0.14y = 970 (2)

Manipulate (1)

x +y = 5000

y = 5000 - x (3)

Substitute (3) into (2)

0.23x + 0.14y = 970

0.23x + 0.14(5000-x) = 970

0.23x + 700 - 0.14x = 970

0.09x = 270

x = 270/0.09 = 3000

He gave away 3000 pens.
Altogether, he gave away 5000 advertisements. Therefore, he gave away 2000 stickers as well.

Hope that helps!